OpenStudy (anonymous):

Problem Posted Below: Could really use the assistance of anyone who is available.

OpenStudy (anonymous):

(A) Make a table of values rounded to two decimal places for the function \[f(x)=e ^{x}\] for x=1,1.5,2,2.5,3. Then use the table to answer parts (B) and (C)  (B) Find the average rate of change of f(x) between x=1 and x=3. (C) Use average rates of change to approximate the instantaneous rate of change of f(x) at x=2.

OpenStudy (anonymous):

So i got Part A done with and im onto Part B, I believe i use the formula:\[\frac{ f(x+h)-f(x) }{ h }\] but not entirely understanding how to use this equation.

OpenStudy (anonymous):

so i looked at it a bit more and think im on the right track, i got my formula to equal \[\frac{ e ^{(x+h)}-e ^{(x)} }{ h }\] and i inserted x=1 and x=3 into the equation getting:\[f(1)=\frac{ e ^{(1+h)} -e ^{1}}{ h }\] and \[f(3)=\frac{ e ^{(3+h)}-e ^{(3)} }{ h }\] but not sure what im supposed to do from here

OpenStudy (anonymous):

I found that \(A=\frac{f(3)-f(1)}{3-1}\)

OpenStudy (anonymous):

well thats to find the instantaneous velocity i believe. im trying to find the average rate of change which my book says the formula for that is \[\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

Oh what is \(h\)?

OpenStudy (anonymous):

@cwrw238 do you have time to help me out with this problem?

OpenStudy (anonymous):

im not really sure actually, thats why im stuck i think